For the function f(x) = 2x − 3 x (a) Use the limit definition of the derivative to find f 0 (x). (b) Find the equation of the line tangent to f(x) at x = 3.


1
Expert's answer
2021-09-30T00:35:37-0400

Given that "f(x)=2x-3"

(a) For "y=f(x)" , the derivative of y (with respect to x ) is

 

"\begin{aligned}

&y^{\prime}=\frac{d y}{d x}=\lim _{\Delta x \rightarrow 0} \frac{\Delta y}{\Delta x} \\

&f^{\prime}(x)=\lim _{\Delta x \rightarrow 0} \frac{f(x+\Delta x)-f(x)}{\Delta x} \\

&f^{\prime}(x)=\lim _{h \rightarrow 0} \frac{f(x+h)-f(x)}{h} \\

&f^{\prime}(x)=\lim _{u \rightarrow x} \frac{f(u)-f(x)}{u-x}

\end{aligned}"

"\therefore f'(0)= \begin{aligned}\lim _{h \rightarrow 0} \frac{f(x+h)-f(x)}{h} \\
\end{aligned}"

"\Rightarrow f'(0)=\lim\limits_{h \rightarrow 0} \frac{2(x+h)-3-2x+3}{h} \\
\Rightarrow f'(0)=\lim\limits_{h \rightarrow 0} \frac{2h}{h}\\
\Rightarrow f'(0)=2"


(b) Given that "f(x)=2x-3"

"\therefore f'(x)=2\\
So, f'(3)=2"

At "x=3," we get: "y=6-3=3"

Equation of tangent:

"y-3=2(x-3)\\
\Rightarrow y=2x+3"